### v1

# VILA-M3: Enhancing Vision-Language Models with Medical Expert Knowledge

Abstract
Generalist vision language models (VLMs) have made significant strides in computer vision, but they fall short in specialized fields like healthcare, where expert knowledge is essential. In traditional computer vision tasks, creative or approximate answers may be acceptable, but in healthcare, precision is paramount. Current large multimodal models like Gemini and GPT-4o are insufficient for medical tasks due to their reliance on memorized internet knowledge rather than the nuanced expertise required in healthcare. Meanwhile, existing medical VLMs (e.g. Med-Gemini) often lack expert consultation as part of their design, and many rely on outdated, static datasets that were not created with modern, large deep learning models in mind. VLMs are usually trained in three stages: vision pre-training, vision-language pre-training, and instruction fine-tuning (IFT). IFT has been typically applied using a mixture of generic and healthcare data. In contrast, we propose that for medical VLMs, a fourth stage of specialized IFT is necessary, which focuses on medical data and includes information from domain expert models. Domain expert models developed for medical use are crucial because they are specifically trained for certain clinical tasks, e.g. to detect tumors and classify abnormalities through segmentation and classification, which learn fine-grained features of medical datafeatures that are often too intricate for a VLM to capture effectively especially in radiology. This paper introduces a new framework, VILA-M3, for medical VLMs that utilizes domain knowledge via expert models. Domain expert models developed for medical use were trained to detect tumors and classify abnormalities (segmentation, classification), which learn fine-grained features of the data compared to coarse features that a VLM can typically process. We argue that generic VLM architectures alone are not viable for real-world clinical applications and on-demand usage of domain-specialized expert model knowledge is critical for advancing AI in healthcare. Through our experiments, we show an improved state-of-the-art (SOTA) performance with an average improvement of 9% over the prior SOTA model Med-Gemini and 6% over models trained on the specific tasks. Our approach emphasizes the importance of domain expertise in creating precise, reliable VLMs for medical applications.
1 Introduction
Large language models (LLMs) are at the forefront of AI. Recent works [28, 32] have enabled LLMs to support visual inputs to expand on their use for many vision applications by transforming them into vision language models (VLM). While foundational VLMs developed for computer vision applications have achieved remarkable success in various general-purpose tasks, they struggle to meet the necessary precision required for medical tasks [37, 47, 5]. VLMs trained on general tasks often lack specialized medical domain knowledge to interpret radiological images correctly [15, 5, 39]. VLMs without specific medical training also often miss the subtle visual details that are critical for medical diagnosis. For example, a recent study on breast imaging [9] shows that generic models like GPT-4o (OpenAI) are not suitable for medical use cases. At the same time, there are healthcare-specific large VLMs such as Med-Gemini (1.5 trillion parameters) [58], which are designed without consideration of expert model information and are likely to neglect finer visual features. Current VLMs are well-adapted to connect coarse or dominating visual features with language. However, they often fail to recognize finer visual details (Fig. 3). There are multiple reasons for these deficits. A major reason is the limited availability of comprehensive and clinically relevant public datasets for medical vision language tasks. Most public datasets in medical imaging were released to address specific narrow AI tasks like classification, regression, and segmentation [29]. At the same time, healthcare datasets incorporating natural language are limited to specific tasks such as visual question answering (VQA) and report generation, which do not cover the breadth of clinically relevant tasks.
Over the years, medical imaging has shown successful applications of specialized narrow AIs that solve specific medical imaging tasks, such as semantic segmentation. These “expert models” have already been incorporated into clinical workflows, such as pre-operative planning, patient triage, and more. Many have received regulatory approval (e.g., FDA) and, therefore, incorporating the “knowledge” of existing expert models into the decision-making and generation process of VLMs, specifically in healthcare, should markedly improve the overall performance on clinically relevant tasks.
In this work, we propose a new VLM framework, VILA-M3, that addresses the unique challenges faced by general-purpose VLMs when applied to the medical domain by incorporating the domain-expert knowledge of existing segmentation and classification models. We evaluate our approach on several medical imaging benchmarks and show significant improvements over the state of the art (SOTA), see Fig. 1.
1.1 Key challenges
Next, we summarize the key challenges that VILA-M3 aims to address.
Limitations of Existing Medical Datasets: The majority of publicly available medical datasets used for model training and evaluation are static and not tailored for training for large VLMs [21, 50, 16, 33, 26, 11, 4]. The datasets were originally designed for earlier machine learning approaches, and they do not reflect the complexities required for modern deep learning models such as VLMs. This need has sparked the generation of newer datasets such as [55, 7, 19]. In medicine, the constant need for expert validation and the evolving nature of healthcare data make static datasets insufficient for accurate model evaluation. In this work, we outline the need for more dynamic and adaptive approaches and show that on-demand usage of domain expert models can lead to improved model performance. The dataset limitation is one of the primary reasons we require and utilize expert model information in our framework. While large models like GPT-4o have been suggested as potential tools for evaluating model performance in medicine, their reliance on memorized information makes them unreliable for precision-critical tasks [15, 5].
Combining General Vision-Language Data with Healthcare Data: Healthcare-specific datasets, while valuable, are not sufficient to fully train models that also require linguistic understanding. To overcome the aforementioned challenge, we utilize a four-stage training schema: pre-training vision encoder, pre-training VLM, instruction fine-tuning (IFT), and IFT with domain expert information. The two stages of IFT ensure that models can preserve their language capabilities without sacrificing performance on medical benchmarks [14]. This approach not only enhances their overall performance but also addresses the unique challenges posed by medical language and its precision requirements.
1.2 Contributions
Motivated by the above challenges, we present the following key contributions of our work:
Expert Knowledge Integration: We emphasize the need for integrating expert knowledge into medical VLMs to improve precision.
Comprehensive capability: VILA-M3 is the first medical VLM that can tackle segmentation, classification, report generation, and VQA tasks in one framework.
Expert-guided instruction fine-tuning: We preserve the VLMs language capabilities by introducing expert-guided IFT training on top of the built VLM training schemes.
Hybrid 2D/3D Information Fusion: We introduce effective integration with 2D and 3D expert models, enabling hybrid fusion of domain expert models that provide relevant 3D spatial information to enhance VLMs limited to 2D inputs.
Open-Source Module: We provide an open-source module for data preparation, training, and model evaluation in medical VLM.
2 Related Work
| Model Type | Visual Question Answering | Report Generation | Classification | ||||||
| Datasets | Rad | MIMIC | SLAKE | Path | CXR (Exp.) | CXR (Exp.) | X-ray14 (Exp.) | CheXpert (Exp.) | Tot. Avg. |
| Metric | Acc. | Acc. | Acc. | Acc. | BLEU-4 | ROUGE | F1 | F1 | Tot. Avg. |
| Med-Gemini (1.5T) | 78.8* | 78.6 | 84.8 | 83.3 | 20.5* | 28.3* | 46.7 | 48.3 | 55.7 |
| Task Spfc. SOTA | 84.2 | - | 86.8 | 91.7 | 15.4 | 30.6 | 50.0 | 51.5 | 58.6 |
| VILA-M3 (3B) | 78.2 | 82.4 | 79.8 | 87.9 | 20.2 | 31.7 | 51.3 | 60.8 | 61.5 |
| VILA-M3 (8B) | 84.7 | 82.1 | 82.7 | 91.0 | 21.1 | 32.0 | 48.9 | 61.6 | 63.0 |
| VILA-M3 (13B) | 80.5 | 86.4 | 83.2 | 91.0 | 21.6 | 32.1 | 51.2 | 61.5 | 63.4 |
| VILA-M3 (40B) | 90.4 | 81.9 | 84.0 | 92.7 | 21.6 | 32.2 | 51.3 | 61.0 | 64.3 |
Vision-language models for medicine: VQA tasks were one of the first that combined text and images directly in a single task [3]. These were later extended to FLAN-style visual instruction tuning tasks [52, 56]. General-purpose large multi-modal models such as GPT-4 [1] and Gemini [46] based on transformers [49] have demonstrated great potential towards building intelligent conversational assistants built on combined textual and vision datasets. When applied in the biomedical domain, the models present promising capabilities of understanding the basic concepts. However, handling complex domain-specific tasks with high accuracy and fine granularity remains both challenging and demanding [58, 28, 42, 37]. Pre-training and instruction tuning is important for aligning the representations across modalities and enhancing the reasoning capabilities [32, 58].
Concurrently, there have been multiple methods targeting large VLMs for medicine in different aspects. Med-Gemini [58] boasts a 1.5 Trillion parameter model and includes non-vison tasks like genomics and medical exams. LLaVa-Med [28], based on the popular LLaVa architecture [34], focused on using academic medical datasets, while Med-Flamingo [37] is an extension of Flamingo [2] that introduced a novel mechanism for combining vision and text. BiomedParse allows promptable segmentations to be performed by the VLM itself [60] but does not generalize to more complex tasks like report generation. While promising first steps towards VLMs in medicine, these works highlight the limitations in precision and lack of domain expertise when applied to medical tasks, compared to traditional computer vision applications on narrow tasks such as classification and segmentation, where traditional CNNs often outperform their VLM counterparts.
Medical vision models and expert systems: Many models have been developed to address domain- and task-specific challenges. They achieve state-of-the-art (SOTA) performances in certain areas and often focus on expert knowledge learned from carefully curated datasets. For example, Myronenko et al. [38] created accurate lesion segmentation models for multi-modal brain MRI, He et al. [17] studied 127 common types of human anatomical structures segmentation from tens of thousands of expert annotated CT examples. Several chest X-ray classification models have been developed by Cohen et al. [8] on a wide range of publicly available datasets. Integrating domain knowledge and guided reasoning while maintaining modular flexible system design can be a plausible approach to leverage the previous successes [53, 39].
Evaluation and benchmarking in healthcare AI: Evaluation for vision-language models is a challenging task and is heavily dependent upon the type of task that is being evaluated. While it is straightforward to evaluate close-ended tasks such as classification, segmentation, or object detection as the well-established metrics can be directly used (e.g., F1, Dice, accuracy, etc.), it is not well-established how to evaluate more open-ended tasks where the answers have a wide range for variations. For example, open-ended questions in VQA and report generation require the evaluation of the correctness of entire sentences or paragraphs, which is challenging [12]. Many prior works attempted to evaluate the open-ended weaker VLM output by a stronger VLM. For instance, ChatGPT-4o is being used as a judge to evaluate the answer [34, 28, 54]. While using stronger VLMs as judges might be acceptable for computer vision tasks, this method of evaluation is not satisfactory for healthcare, where a factual level of precision is required, and patient outcomes depend on the results. Our evaluation shows that the capabilities of off-the-shelf VLMs like ChatGPT-4o are too limited to judge the outputs of medical VLMs directly.
Agentic systems: The flexibility of generative AI provides new possibilities for developing powerful medical agents that can integrate heterogeneous information and generic reasoning. Most previous works focus on extending large language models in a medical context [45, 57, 44, 30, 51]. Very recently, Hoopes et al. [18] developed a code-predicting agent framework that can integrate an array of external systems and APIs for a set of neuroimaging tasks including brain region analysis. Li et al. [27] proposed a multi-modal large language model invoking task-specific tools and aggregating results from them.
3 Method
The VILA framework [32] is suitable for pre-training vision and language models as well as instruction fine-tuning (IFT) them with domain-specific datasets. VILA-M3, in particular, leverages domain-expert model insights alongside the versatile capabilities of VILA to capture the nuances of radiological language and the intricate details of medical images, enabling more precise image-text fusion. It enhances tasks such as visual question answering (VQA), zero-shot image classification, image retrieval, and report generation, and it even incorporates expert model feedback to produce more accurate and well-reasoned outputs. VILA is particularly applicable for medical AI, where accurate association between visual data and textual information is critical for clinical decision-making. We further enhance it by including expert model feedback that can be triggered on demand.
3.1 Training Framework
The standard approach to train VLMs like VILA is to use a corpus of visual-text data where text based conversations are associated with an image. The image token is randomly inserted in between the text or before/after the text (Fig.1). As a common practice that the inclusion of text-only data is helpful towards preserving the language capabilities of VLM, we included a small corpus of medical text data [22] to specifically prevent degradation of the LLM capabilities as the model gets conditioned towards medical images.
VILA is based on auto-regressive multi-modal LLMs, where images are tokenized into visual tokens, concatenated with text tokens, and fed into language models, effectively treating visual input as a foreign language (Fig.1). Therefore, a projector layer is needed to bridge the token embeddings from the image and text modalities. While different projector architectures are possible, a simple linear layer encourages the LLM to generalize better [32]. This approach naturally extends text-only LLMs by incorporating visual embeddings, allowing them to handle interleaved image-text inputs and generate text outputs. There are other frameworks available that combine vision and language, such as cross-attention based [37, 2], but auto-regressive approaches are gaining in popularity due to their flexibility in supporting different vision encoder and LLM backbones. Therefore, we chose VILA as it is the current SOTA for traditional computer vision and is an open-source framework. VILA models are typically trained in three training stages: projector pre-training, visual language pre-training, and IFT. VILA-M3 builds on the last stage of visual instruction tuning and expands it with the inclusion of information from domain expert models. While VILA has enough flexibility to allow for separate fine-tuning of the vision encoder, the projector, and the LLM, we fine-tune all as it is necessary to customize both vision, projector alignment, and LLM for the medical domain as shown in Fig. 2.
3.2 Expert Models Triggering & Feedback
In this work, we use open-source expert models for common medical imaging tasks. VILA-M3 learns to predict a keyword and argument, e.g., <VISTA3D(hepatic tumor)> when it wants to trigger a suitable expert model. The list of available expert models is fed as a system prompt to the model (see Fig. 2).
Used Expert Models: For volumetric segmentation in CT images, we utilize VISTA3D [17]. VISTA3D is a domain-specialized interactive foundation model developed for segmenting and annotating human anatomies with precision. It achieves a highly accurate segmentation of 127 classes, including challenging tumors. We include several organ or anatomy subgroups that can be chosen by VILA-M3 using different arguments when predicting the model keyword. Valid arguments are ‘everything’ (i.e., all 127 available organ classes), ‘hepatic tumor’, ‘pancreatic tumor’, ‘lung tumor’, etc. These arguments are described in model cards available to VILA-M3 as context at training and inference time (see Fig. 1). Note that when triggered, VISTA3D can segment the full volumetric scan even though VILA-M3 only processes 2D inputs.
For MRI imaging, we focus on tumor segmentation in multimodal brain MRI. The MONAI BRATS model [38] consists of a pre-trained model for volumetric (3D) segmentation of brain tumor sub-regions from four MRI modalities (T1-weighted, T1-weighted with gadolinium contrast-enhancement, T2-weighted, and FLAIR) based on BraTS 2018 data and achieves good performance [36].
The majority of the data sets used for training contain chest X-rays (CXRs), which are used for many routine diagnostic clinical tasks. In this work, we utilize a model ensemble of CNN-based classification models from TorchXRayVision [8]. These models are trained on large cohorts of data and provide good accuracy across several diverse CXR datasets.
Expert Model Feedback: When VILA-M3 selects and activates an appropriate expert model from the available model cards, the result is returned to VILA-M3 as another user prompt, formatted in a conversational style. For a segmentation result, we re-tokenize the generated mask and provide a textual description that explains the significance of the segmentation results, overlaid on the original image(see Fig. 3). In the case of classification, we fed back the result as a list of yes/no statements for the likelihood that the image contains a certain disease (i.e., a list of 18 diseases using the TorchXRayVision ensemble).
3.3 Datasets
Multiple publicly available datasets were utilized for this study. We present a brief overview of them and the processing that we applied to effectively utilize them for training with the VILA framework. All training data is detailed in Supp. material.
VQA: In this study, we utilize the RadVQA, SLAKE, and PathVQA datasets to evaluate visual question answering (VQA) performance in medical imaging [16, 33, 26]. The SLAKE dataset comprises 642 medical images across modalities such as CXR, CT, and MRI, accompanied by 14,028 question-answer pairs covering radiology across various anatomical regions. From the official dataset splits, we used 4,919 training, 1,053 validation, and 1,061 test examples, focusing on diverse question formats like open-ended and yes/no. The PathVQA dataset includes 4,289 pathology images paired with 32,632 question-answer pairs, divided into 19,654 for training, 6,259 for validation, and 6,719 for testing. Both datasets probe a wide range of visual and clinical attributes, including abnormalities, organ identification, and diagnostic information. Our analysis focuses on using these datasets to enhance the performance of medical vision language models. We utilized both the training and validation sets from all the VQA datasets as part of the model training, due to the huge compute cost of training VLM’s standard validation techniques are quite compute heavy and also redundant.
Classification: We adopted two publicly available high-quality datasets of abnormalities in CXR images as a benchmark for the model’s visual classification capability. One is the subset of ChestX-ray14 [50] curated by Majkowska and Mittal et al. [35], consisted of 1,962 images after an adjudication process of four radiologists. For this dataset, we follow [58] and focus on the classification of three conditions – lung opacity, pneumothorax, and fracture. The other dataset is a subset of CheXpert [21] created by Saporta et al. [43] on the labeling of five pathological conditions, namely atelectasis, cardiomegaly, consolidation, edema, and pleural effusion.
Report Generation: For report generation, we adopt the MIMIC-CXR111v2.1.0 dataset [23, 24], which provides over 377,110 CXR images in JPG format, along with structured labels extracted from 227,827 associated free-text radiology reports. Each image is linked to the corresponding radiology report, which contains the radiologists’ findings and impressions. Labels for the dataset were derived using automated tools CheXpert [21] and NegBio [41], which extract and classify medical conditions (e.g., “pneumonia,” “edema,” “pleural effusion”) as positive, negative, uncertain, or not mentioned based on the language of the reports. Additionally, radiologists manually annotate a portion of the report, providing 14 categories of findings for model evaluation and validation. The dataset is divided into training, validation, and testing parts, each with separate CSV files to support reproducible research. Due to suboptimal report quality in the initial datasets, we applied advanced Large Language Models (LLMs) to refine text reports through targeted noise reduction and finding extraction. This process significantly improved model accuracy and reliability. Additional methodological details are available in the supplementary materials.
Expert Datasets: To create training data that triggers expert models for tasks like segmentation and classification, we leverage existing datasets by running expert model inference. For instance, we can segment existing CT datasets with the VISTA3D model, using the inference results to build training conversations tailored for VILA-M3. Similarly, with a brain MRI dataset, we can run the BRATS segmentation model and then utilize the output to generate structured training conversations for training VILA-M3. This approach facilitates the generation of context-specific dialogues that enhance model alignment with expert-level tasks for medical imaging. Further, we extend a subset of the existing VQA and report generation tasks by adding the model cards to the message context and adding trigger statements for the corresponding CXR expert model as illustrated in Fig. 1.
4 Experiments & Results
4.1 Model Architecture Hyper-parameters
We use the image resolution in the visual encoder at 384384 using OpenAI’s CLIP-L, enabling the model to capture more fine-grained visual details. For the LLM backbones, we use Vicuna (based on Llama-2 [48]) for 3 Billion and 13 Billion parameter checkpoints, for 8 Billion, we utilize Llama-3 [10] as the backbone and the 40 Billion uses Yi-34B as the LLM backbone based on Yi [59] and uses a 448448 vision encoder. The LLM choices depend on the configurations used for the VILA checkpoints pre-trained on natural images.
4.2 Implementation Details
Training Hyper-parameters: To ensure a fair comparison for all VILA model size variants, we kept the same training hyper-parameters for each size. A consistent batch size of 16 was maintained at the device (per GPU) level for training and 4 for evaluation. Gradient accumulation was adjusted accordingly to ensure the same overall effective batch size in multi-node training. The learning rate was initialized at 2e-5, using a cosine learning rate schedule with a warm-up ratio of 0.03. Weight decay was set to zero to avoid over-regularization, and all the models utilized bf16 mixed precision, allowing efficient training with a reduced memory footprint. Gradient check-pointing was enabled to further optimize memory usage. The LLM backbone, projector, and vision encoder were kept unfrozen as that is the best setting to adapt for new data as per [32] to fine-tune the entire model.
Training Cost: To train these models we used A100’s with 80GB of memory. For the 3, 8, and 13 billion parameter models, we utilized four nodes (x8 A100) in parallel using 32 GPUs in total. For the 40B model, we utilized sixteen nodes, utilizing a total of 128 GPUs in parallel (Table 2). More information regarding inference computational cost can be found in the supplementary material.
Inference Cost: With the open source release222github url [BLINDED] the computational footprints are as follows: The CXR expert dynamically loads various TorchXrayVision models and performs ensemble predictions. The memory requirement is circa 1.5GB in total. The VISTA3D expert model dynamically loads the VISTA3D model to segment a 3D-CT volume. The memory requirement is roughly 12GB, and peak memory usage can be higher, depending on the input size of the 3D volume.
| #Params | Nodes | Tot. GPUs | Epochs | Training Time |
| VILA-M3-3B | 4 | 32 | 2 | 5.5 hrs |
| VILA-M3-8B | 4 | 32 | 2 | 11.0 hrs |
| VILA-M3-13B | 4 | 32 | 2 | 19.5 hrs |
| VILA-M3-40B | 16 | 128 | 2 | 21 hrs |
| Model | LLM | Params | VQAv2 | GQA | VizWiz | SQA-I | VQA-T | POPE | MME | MMB | MMB-CN | SEED | MM-Vet | Avg |
| VILA1.5 | Sheared LLaMa | 3B | 80.4 | 61.5 | 53.5 | 69.0 | 60.4 | 85.9 | 1442.4 | 63.4 | 52.7 | 60.9 | 35.4 | 63.2 |
| VILA-M3 | Sheared LLaMa | 3B | 73.9 | 56.0 | 38.4 | 49.5 | 53.3 | 84.0 | 1381.0 | 60.6 | 51.5 | 59.2 | 23.5 | 56.2 |
| VILA1.5 | Llama-3 | 8B | 83.0 | 63.5 | 63.2 | 82.0 | 68.5 | 85.6 | 1634.9 | 75.3 | 69.9 | 66.4 | 43.2 | 71.1 |
| VILA-M3 | Llama-3 | 8B | 70.4 | 53.2 | 38.4 | 72.3 | 51.2 | 80.8 | 1530.0 | 68.8 | 63.3 | 63.1 | 28.2 | 60.5 |
| VILA1.5 | Vicuna-V1 | 13B | 82.8 | 64.3 | 62.6 | 80.1 | 65.0 | 86.3 | 1569.5 | 74.9 | 66.3 | 65.1 | 44.3 | 70.0 |
| VILA-M3 | Vicuna-V1 | 13B | 78.4 | 61.2 | 55.9 | 74.1 | 60.2 | 83.7 | 1531.0 | 71.3 | 65.5 | 64.8 | 36.6 | 66.2 |
| VILA1.5 | Yi-34B | 40B | 84.3 | 64.6 | 62.2 | 87.2 | 73.6 | 87.3 | 1726.8 | 82.4 | 80.2 | 69.1 | 53.0 | 75.5 |
| VILA-M3 | Yi-34B | 40B | 70.2 | 50.7 | 41.3 | 0.0 | 53.0 | 82.9 | 1548.5 | 58.9 | 55.6 | 56.8 | 27.4 | 52.2 |
| Model | MIMIC-CXR w/o Expert | MIMIC-CXR with Expert | |||||
| Type | Params | BLEU-4 (↑) | ROUGE (↑) | GREEN (↑) | BLEU-4 (↑) | ROUGE (↑) | GREEN (↑) |
| Med-Gemini [58] | 1.5T | 20.5 | 28.3 | - | 20.5 | 28.3 | - |
| Llava-Med [28] | 7B | 1.0 | 13.3 | - | 1.0 | 13.3 | - |
| Llava-Rad [6] | 7B | 15.4 | 30.6 | - | 15.4 | 30.6 | - |
| DCL [31] | 0.25B | 10.9 | 28.4 | - | 10.9 | 28.4 | - |
| VILA-M3-3B | 3B | 19.7 | 31.4 | 39.4 | 20.2 | 31.7 | 39.1 |
| VILA-M3-8B | 8B | 21.4 | 32.2 | 39.8 | 21.2 | 32.1 | 39.7 |
| VILA-M3-13B | 13B | 21.4 | 32.1 | 39.3 | 21.6 | 32.1 | 39.3 |
| VILA-M3-40B | 40B | 21.6 | 32.2 | 39.2 | 21.6 | 32.2 | 39.2 |
4.3 Benchmarking
Evaluation Metrics: Accuracy as a measure was used to gauge the performance of close-ended questions for VQA tasks. F1-score was utilized for classification tasks. For report generation, we utilize three different metrics: BLEU-4, ROUGE, and GREEN score, which aims to be a more clinically relevant measure [40]. In general, report generation is more challenging to evaluate and the metrics present different perspectives to understand the performance in a comprehensive manner.
VILA-M3 Outperforms SOTA: As a baseline comparison, we use the results from the Med-Gemini work directly [58] and task-specific SOTA baselines [28, 31, 6, 8]. Overall, the results observed in Fig. 1 indicate superior performance to existing SOTA methods that are available from either open-source or closed-source. In particular, it can be seen that markedly higher performance can be achieved in 7 out of 8 metrics with much smaller models (a few billion parameters) compared to Med-Gemini, which contains 1.5 trillion parameters on this medical benchmark.
Three models per parameter size variant were trained for 1, 2, and 3 epochs, respectively. The best performing VILA-M3 model was determined via comparison of all parameter size variants observed in Table. 1.
VILA-M3 Benchmark: We also evaluated VILA-M3 models on the VILA benchmark to ensure that they are not overfitting and degrading their baseline capabilities. Table. 3 summarizes the results on the base VILA checkpoint and the VILA-M3 checkpoints after expert-guided IFT.
4.4 Effectiveness of Expert-guided IFT
Three key results prove the effectiveness of expert-guided IFT. The first is the observation that VILA-M3 outperforms SOTA models (Fig.1 & Table.1). The second major observation is that the degradation on the VILA benchmark is not significant, after extensive training with expert-guided IFT, the performance of VILA-M3 only degrades by 7%, 11% and 4% for the 3B, 8B and 13B model variants (Table. 3). Contrastingly, the 40B model degrades by 23%, we attribute this to Llama models being superior than Yi [59] models. The third observation can be drawn that VILA-M3 is not over-fitting to the expert IFT training process for epoch 1 and 2 models. This can be visually observed in Fig. 4 that over-fitting happens for epoch 3 models, these results are shown for the 8B model (results for more model variants are available in supplementary material).
| Model | ChestX-ray14 Classif. w/o Expert | ChestX-ray14 Classif. With Expert | Average | ||||||
| Type | # Params | Fracture | Pneumothorax | Lung opacity | Fracture | Pneumothorax | Lung opacity | w/o Expert | Expert |
| GPT-4o | Unknown | 2.0 | 19.7 | 77.8 | - | - | - | 33.1 | - |
| Med-Gemini | 1.5T | 5.5 | 55.3 | 79.9 | - | - | - | 46.9 | - |
| TorchXrayVision | (ensemble) | - | - | - | 11.6 | 50.4 | 87.9 | - | 50.0 |
| VILA-M3-3B | 3B | 7.0 | 54.5 | 83.8 | 10.8 | 55.2 | 88.1 | 48.4 | 51.3 |
| VILA-M3-8B | 8B | 2.7 | 49.9 | 85.0 | 14.4 | 53.5 | 84.3 | 45.8 | 50.8 |
| VILA-M3-13B | 13B | 7.6 | 55.9 | 86.2 | 13.6 | 51.5 | 88.5 | 49.9 | 51.2 |
| VILA-M3-40B | 40B | 0.0 | 55.9 | 85.3 | 11.0 | 53.9 | 89.0 | 47.1 | 51.3 |
| Model | CheXpert Classif. w/o Expert | CheXpert Classif. With Expert | Average | ||||||||||
| Type | # Params | Atel. | Cardio. | Consol. | Edema | Pl.Eff. | Atel. | Cardio. | Consol. | Edema | Pl.Eff. | w/o Expert | Expert |
| GPT-4o | Unknown | 31.8 | 49.6 | 13.8 | 37.2 | 36.8 | - | - | - | - | - | 33.9 | - |
| Med-Gemini | 1.5T | 49.7 | 72.0 | 23.0 | 32.7 | 64.4 | - | - | - | - | - | 48.4 | - |
| TorchXrayVision | (ensemble) | - | - | - | - | - | 61.5 | 63.8 | 28.6 | 50.5 | 52.9 | - | 51.5 |
| VILA-M3-3B | 3B | 59.0 | 69.3 | 24.0 | 71.8 | 62.9 | 61.2 | 71.5 | 35.0 | 63.7 | 72.4 | 57.4 | 60.7 |
| VILA-M3-8B | 8B | 63.5 | 63.2 | 35.1 | 73.1 | 71.9 | 62.7 | 70.6 | 36.7 | 68.0 | 64.5 | 61.4 | 60.5 |
| VILA-M3-13B | 13B | 61.9 | 71.1 | 20.5 | 65.4 | 60.2 | 62.8 | 76.0 | 39.4 | 67.1 | 62.5 | 55.8 | 61.5 |
| VILA-M3-40B | 40B | 57.2 | 70.4 | 32.6 | 68.3 | 68.7 | 63.4 | 72.1 | 38.4 | 63.7 | 67.4 | 59.4 | 61.0 |
4.5 Ablation Studies
Expert Model Effect: To validate the findings of domain-expert models, we conduct experiments with VILA-M3 using the same training setting as outlined above in section 4.2. In particular, for the classification with and without expert feedback, we use the baselines of TorchXrayVision [8] and also compare it with the flagship model GPT-4o [1, 20] from OpenAI for both datasets ChestX-ray-14 and CheXpert, see Tables 5 and 6.
Experiments were performed for report generation with and without expert information inclusion (Table 4). For comparison, we also utilize established baselines for the task-specific state-of-the-art (SOTA) “Dynamic Graph Enhanced Contrastive Learning” (DCL) model [31]. Notably, other baseline models considered in the study included Llama-Med [28], Llama-Rad [6], and Med-Gemini [58], in addition to the DCL models [31].
For both classification and report generation tasks, the improvements achieved by inclusion of expert models can be observed in Tables 4, 5 & 6. It can also be observed that large dataset models such as GPT-4o do not perform well on these tasks.
For segmentation tasks, observing Fig. 3 reveals that our model VILA-M3 fails to pick up on tumor findings if segmentation is not provided. However, with additional expert information included it is able to detect the finer features of a tumor.
Balanced vs. Unbalanced Datasets: The original size of datasets varies a lot, especially when the original numbers are compared between VQA datasets and MIMIC-CXR. Therefore, we increased the frequency of low-count datasets in a category-wise manner (based on VQA, expert, report gen., and classification type) to systematically balance the training dataset. The training settings were kept the same as outlined in section 4.2. The training dataset table with frequency counts can be found in the supplementary materials.
The improvements based on the balanced training dataset as compared to the unbalanced dataset (original dataset size) can be observed in Fig. 5. Quantitatively, an average improvement of 4% of all metrics is gained by balancing.
4.6 Training Convergence
Despite different model architectures and sizes, we observe that the proposed training procedures converge successfully in all cases and are not sensitive to the choice of hyper-parameters. Taking the models trained with 4 nodes (32 GPUs) as an example, Fig. 6 shows that given the same training dataset, the training loss scales with model size as expected according to the empirical power law [25].
5 Discussion & Conclusion
The proposed VILA-M3 model represents a significant advancement in multimodal machine learning for healthcare. It achieves SOTA performance as both a generalist model and one incorporating expert model responses to solve complex tasks. This capability allows the model to maintain high levels of performance when addressing general tasks while also leveraging expert knowledge to improve accuracy. With additional expert models, it can be seamlessly extended to various capabilities, such as registration, regression, etc. Our work can be seen as a step towards enabling “chain of thought” capabilities in VLMs by letting the model reason about when to trigger an expert and how to incorporate its result in the final generated prediction. For future work, we will explore the integration of Retrieval-Augmented Generation (RAG) to further enhance VILA-M3 by retrieving and incorporating relevant information from large datasets dynamically during inference. Additionally, we aim to expand VILA-M3 towards a multi-agent framework that directly incorporates smart expert models that can themselves decide whether to involve additional experts for consultation [13] to optimize performance across varied tasks.
In conclusion, the proposed VILA-M3 framework can be used to train SOTA models that consider expert model responses to improve performance on particular tasks while simultaneously exhibiting good generalist performance.